OpenStudy (anonymous):

The world population in the year 2011 was approximately 7 x 109 people. If the population doubles in 100 years, how many people will be on the earth in the year 2111? Write your answer in scientific notation.

OpenStudy (anonymous):

1526

OpenStudy (anonymous):

\[7\times 10^9\] right?

OpenStudy (anonymous):

double that number to get \(14\times 10^9=1.4\times 10^{10}\)

OpenStudy (anonymous):

in scientific notation 1.526*10^3

OpenStudy (anonymous):

\[(7*109)^{2^{n}}\] n=(2111-2011)/100

OpenStudy (anonymous):

Look at the number line I_________I_____________I__________________I........______________I 2011 2111 YEAR To get from the year 2011 to 2111 you have (2111-2011)=100 Years You stated that the population doubles in 100 years. What does that mean algebraically? Well if you have x and it doubles you have 2x, so you multiply by 2. So if you started with a population say 5.3X10^7 and it doubles you basically do the following: 2*(5.3X10^7) 2*5.3 X 10^7 10.6 X 10^7 But recall scientific notation we need that 10.6 to be less than 10. We simply move the decimal to the left over one space 1.06 X 10^ ? What goes in that question mark? So 10.6 X 10^7= 106000000. you are moving the decimal 7 places over If you want the same number but expressed correctly in scientific notation you have 1.06 X10^8 because this equals 1060000000 since it took one more space to the right to move over to get 106000000.